OpenStudy (anonymous):

What's the answer for this ? -5x+11=1/2y

OpenStudy (texaschic101):

what are you solving for ?

OpenStudy (anonymous):

Write in standard form with interger coefficents?

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):



OpenStudy (texaschic101):

standard form : Ax + By = C -5x + 11 = 1/2y ---- subtract 11 -5x = 1/2y - 11 --- subtract 1/2y -5x - 1/2y = -11 -- standard form has no fractions, multiply by 2 -10x - y = -22 -- multiply by -1 to make x positive 10x + y = 22

OpenStudy (anonymous):

Thank You! can i ask something else?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

Write the standard form of an equation of the line passing through (-5,1) with slope 3/4?

OpenStudy (texaschic101):

y = mx + b 1 = 3/4(-5) + b 1 = -15/4 + b 1 + 15/4 = b 4/4 + 15/4 = b 19/4 = b your equation is y = 3/4x + 19/4 but we need Ax + By = C so... y = 3/4x + 19/4 -- subtract 3/4x -3/4x + y = 19/4 -- no fractions, multiply by 4 -3x + 4y = 19 -- multiply by -1 to make x positive 3x - 4y = -19 -- standard form

OpenStudy (anonymous):

So wsrite both of them down? or jus the first one?

OpenStudy (texaschic101):

if you need to show how you got the answer, you might need to write both

OpenStudy (texaschic101):

if you just need the answer, write 3x - 4y = -19

OpenStudy (anonymous):

Ok.. Write the standard form of the equation? a. a horizontal line thorugh (6,-5) b. vertical line thorugh (-2,7)